the sum of two numbers is 63. the first number is 2/5 of the second number. what are the numbers?
Answer by DrBeeee(684)

Let a = the first number
Let b = the second number
From the problem statement we have
(1) a = 2/5b and
(2) a + b = 63
Put a of (1) into (2) and get
(3) 2/5b + b = 63 or
(4) 7/5b = 63 or
(5) b = (5/7)*63 or
(6) b = 5*9 or
(7) b = 45
Then from (2) we get
(8) a + 45 = 63 or
(9) a = 63 - 45 or
(10) a = 18
Check this with (1).
Is (18 = (2/5)*45)?
Is (18 = 2*9)?
Is (18 = 18)? Yes
Answer: The two numbers are 18 and 45.